OpenStudy (lina17):

4(2x-1)-3=9(x-4) help?

OpenStudy (priyar):

POINTS: 1.An equation implies both sides of the "equal to" sign are the same. 2.X is an unknown quantity(A quantity =number(representative of some physical quantity)). 3.A simple bracket means that you need to multiply the terms(all of them) by the number outside the bracket.

OpenStudy (priyar):

where exactly are you stuck ?

OpenStudy (priyar):

ok @lina17.. what do u get after multiplying 4 inside...

OpenStudy (lina17):

(8x-4)-3=9(x-4)

OpenStudy (priyar):

that is right

OpenStudy (priyar):

simplification: This means you are trying to simplify your problem by reducing the number of terms.You can do so by adding like terms,multiplying inside the bracket,etc....ON BOTH SIDES OF THE EQUATION.....

OpenStudy (lina17):

so i multiply 9(x-4)

OpenStudy (priyar):

yes

OpenStudy (lina17):

so it would be (8x-4)-3=9x-36

OpenStudy (priyar):

that is correct.good job! now if you have nothing to do with the terms inside a couple of brackets,then it is ok to remove them...

OpenStudy (lina17):

so do i subtract 8x-4

OpenStudy (priyar):

8x and 4 are subtractable once you know what x is.... but right now you cant subtract them...there is no way to do so......however ,look at the -4 and the -3 ...they are both similar,as in they are subtractable(that is why they are like terms).

OpenStudy (lina17):

oh okay

OpenStudy (lina17):

so it would be 8x-1=9x-36

OpenStudy (priyar):

remember you are subtracting 3 from negative four not positive four......

OpenStudy (lina17):

so its 8x--1=9x-36

OpenStudy (priyar):

lets go over this: (-4) - (3) = (-4) + (-3) =?

OpenStudy (lina17):

-7 ?

OpenStudy (priyar):

kudos!

OpenStudy (priyar):

so what will be the answer?

OpenStudy (lina17):

so its 8x-7=9x-36

OpenStudy (priyar):

yup

OpenStudy (lina17):

do i subtract 9x-36

OpenStudy (priyar):

try and think of what you will get if you were able to subtract them....is it possible to?

OpenStudy (lina17):

no

OpenStudy (lina17):

this is the part im stuck on

OpenStudy (priyar):

good... so bring the terms which you can add/sub/mul/div together.....

OpenStudy (priyar):

forget about the equal to sign for a minute ........just draw columns and write "like terms" in one column.........for eg: 8x and 9x are like terms put them in one column... similarly,tell me how many classifications/groups of terms are there?

OpenStudy (lina17):

2

OpenStudy (lina17):

?

OpenStudy (priyar):

you are right.......one is the constants and the other is the x terms

OpenStudy (priyar):

now focus on the equation and bring like terms together by moving them from one side of the equation to the other

OpenStudy (lina17):

so do i multiply ???

OpenStudy (priyar):

what do you intend to multiply?

OpenStudy (lina17):

8x and 9x

OpenStudy (lina17):

or no ??

OpenStudy (priyar):

no

OpenStudy (lina17):

so what do i do

OpenStudy (priyar):

imagine a situation where you have four types of goods to sell

OpenStudy (lina17):

ok

OpenStudy (lina17):

so is the answer 29

OpenStudy (lina17):

???

OpenStudy (priyar):

you have four bags: chocolates flour water sugar ........now someone steals some chocolate.....so in your inventory you will subtract it.. FROM THE TOTAL NUMBER OF CHOCOLATES THAT YOU HAVE.....NOT FROM FLOUR......OR ANY OF THE OTHER BAGS.....that is what you mean when you say.....'you can perform arithmetic operations only on like terms'..because it would not make sense any other way. And in case you are wondering why a constant can be multiplied to x terms inside of a bracket ....it just means.....you have multiplied the no. of goods you had...you can double or triple your bags of chocolates......that makes physical sense..right?

OpenStudy (lina17):

because i subtracted 8x-9x = 1x which is just x and then i subtracted -36-7 =29

OpenStudy (lina17):

??

OpenStudy (priyar):

that step is wrong

OpenStudy (priyar):

did you understand my analogy for why we can perform operations on like terms?

OpenStudy (lina17):

kind of

OpenStudy (priyar):

i suggest you review your number line,basic addition and subtraction with negative numbers etc....in this case,8x-9x,you are subtracting a bigger no from a smaller no.

OpenStudy (priyar):

subtract while you count on your fingers......8 =8. so till 8 you can bend a finger for each number.after 8 you have no fingers left to bend down...so you go to the other side of zero giving you -1

OpenStudy (priyar):

??

OpenStudy (priyar):

@lina17 ........now tell me ..8x-9x=?

OpenStudy (priyar):

hope you understood........

OpenStudy (anonymous):